Ask
your own question, for FREE!
Mathematics
11 Online
The probability that rose will complete a free through basketball shot is 1/3. If she makes 5 shots, what is the probability of her completing at least one shot?
Still Need Help?
Join the QuestionCove community and study together with friends!
this is a binomial problem. we want completing at least one shot. Let X = # of shots in 5 chances . Then X can equal to 0,1,2,3,4,5 We want P( X >= 1 ) or P (X = 1,2,3,4,5)
each shot has probability of success 1/3 and failure 2/3 . it is easier to do the complement P ( X > = 1 ) = 1 - P(X <1 ) = 1 - P( X = 0 ) = 1 - ( 2/3)^5
.8683127572
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
9 hours ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal